# Physics and Deep Learning in Computational Wave Imaging

### v1

@@ IV-B2 Physics-Aware Methods
This group of methods incorporates physical knowledge, such as the wave equation, through the use of explicit regularization. Several papers have demonstrated the benefits of using the wave equation to regularize the inversion [74, 11, 75, 60, 53, 76]. Particularly, Gupta [60] proposes to jointly solve both forward and imaging problems by leveraging an invertible neural network as shown in Fig. 4b. The idea is to leverage an invertible U-net to learn invertibility in the latent space where the input and output dimensions are strictly equal. Encoder and decoder blocks are added on both sides of the input and target domains to learn meaningful latent representations. Then, the input data are passed to a latent encoder, and the outputs of this encoder are fed to the invertible U-net. Finally, the outputs from the invertible network are decoded to make a prediction in the target domain. As demonstrated by Gupta [60], this network architecture comes with a major benefit in that it provides flexibility in learning both forward and inverse transformations with a single architecture and further constrains each other during the training, thus the network provides both enhanced imaging and forward modeling predictions.
Incorporating the governing wave equation into ML models has effectively enhanced their performance. Various network structures and parameterization schedules have been explored. Ren et al. [74] develop SWINet, a waveform inversion network that utilizes wave equation-based cells, treating the velocity map as a trainable parameter. Similarly, Yang and Ma [76] introduce FWIGAN, which merges the wave equation with a Wasserstein GAN framework and also considers velocity as trainable. Rasht-Behesht et al. [75] propose a method using a physics-informed neural network [77] to solve wave propagation and constrained wave inversion simultaneously, producing the velocity map from the inversion network. Key challenges for all those methods include the need for an accurate initial guess to start the process, which can be tricky, and the substantial computational demands of solving the wave equation, particularly for three-dimensional problems.
@@ IV-B3 Parameterization Solution Methods
Physics knowledge can be represented by the parameter space of the neural networks. Ulyanov et al. [58] develop the deep image prior, a type of convolutional neural network, to learn useful prior knowledge via re-parameterization. A similar idea has been developed to solve CWI problems as shown in Fig. 4c. In particular, velocity maps are parameterized while the forward model is used to calculate the associated waveform and obtain the data misfit. Mathematically, parameterization can be posed as
$\widehat{\bm{m}}$ $=g_{\bm{\theta}^{*}(\bm{d})}(\bm{\beta}),\;\mathrm{~{}~{}~{}~{}~{}~{}~{}s.t.}\;\bm{\theta}^{*}(\bm{d})=\operatorname*{arg\,min}_{\bm{\theta}}\mathcal{L}(f(g_{\bm{\theta}}(\bm{\beta})),\bm{d})\,,$ (10)
where $\bm{\beta}$ is a random tensor drawn from a normal distribution; i.e., $\bm{\beta}\sim\mathcal{N}(0,\sigma^{2})$ and $g_{\bm{\theta}^{*}}(\cdot)$ is a network model used to parameterize the velocity maps.
@@
Parameterization solution methods, building on conventional physics-based approaches, inherit similar challenges. Typically, these methods require a well-defined initial velocity map, and factors like regularization and optimization techniques (e.g., gradients and step sizes) significantly influence their convergence. Specifically, crucial to Eq. (10) are the selection of $\bm{\beta}$ and the network structure as in $g_{\bm{\theta}^{*}}(\cdot)$. Early work by Wu and McMechan [78] samples $\bm{\beta}$ from noise and used CNNs. He and Wang [57] adopt a similar approach but with a deeper theoretical focus. Both methods involve a two-step process: network pre-training with a specific initial guess, followed by fine-tuning. The discretization assumption limits the output to fixed sizes, restricting transferability. Overcoming these drawbacks, Zhu et al. [56] develop a new strategy that integrates the wave equation for regularization, eliminating the pre-training step and addressing the ill-posed nature of CWI problems. More recently, Sun et al. [55] introduce an approach that employs a continuous implicit function via neural networks to represent velocity maps, accommodating arbitrary sizes.
IV-B4 Plug and Play Methods
The Plug and Play Priors (PPP) method [79, 80, 81], which is widely used in computational imaging, was originally proposed as an approach to exploiting state-of-the-art image denoisers such as BM3D [82] to solve more general inverse problems, but now also represents a convenient approach to exploiting a learned model, such as the DnCNN denoiser [83], without the need to train the model for the specific inverse problem of interest.
It is based on proximal optimization algorithms [84], such as the Alternating Direction Method of Multipliers (ADMM) [85, 86, 87], a fundamental component of which is the scaled proximal operator of a function $h({\bm{m}})$ [88, Ch. 6], defined as
$\operatorname{{prox}}_{\alpha h}(\bm{u})=\operatorname*{arg\,min}_{\bm{m}}\;\alpha h(\bm{m})+\frac{1}{2}\|\bm{m}-\bm{u}\|_{2}^{2}\;.$ (11)
The ADMM algorithm for solving a problem of the form $\operatorname*{arg\,min}_{\bm{m}}h(\bm{m})+r(\bm{m})$, where $h(\cdot)$ is typically a data-fitting term such as $h(\bm{m})=\left\|\bm{d}-f(\bm{m})\right\|_{2}^{2}$, and $r(\cdot)$ typically represents a prior/regularization term, consists of the iterations
$\bm{m}^{(k+1)}$ $=\operatorname*{arg\,min}_{\bm{m}}\;h(\bm{m})+\frac{\rho}{2}\|\bm{m}-\bm{y}^{(k)}+\bm{u}^{(k)}\|_{2}^{2}$ (12)
$=\operatorname{{prox}}_{\rho^{-1}h}\big{(}\bm{y}^{(k)}-\bm{u}^{(k)}\big{)}$
$\bm{y}^{(k+1)}$ $=\operatorname*{arg\,min}_{\bm{y}}\;r(\bm{y})+\frac{\rho}{2}\|\bm{m}^{(k+1)}-\bm{y}+\bm{u}^{(k)}\|_{2}^{2}$
$=\operatorname{{prox}}_{\rho^{-1}r}\big{(}\bm{x}^{(k+1)}+\bm{u}^{(k)}\big{)}$
$\bm{u}^{(k+1)}$ $=\bm{u}^{(k)}+\bm{m}^{(k+1)}-\bm{y}^{(k+1)}\;,$
where $\rho$ is an algorithm parameter that determines the rate of convergence, $\bm{y}$ is an auxiliary variable, and $\bm{u}$ is the Lagrange multiplier associated with the constraint $\bm{y}=\bm{m}$. The PPP method was inspired by the observation that the proximal operator has the form of a maximum a posteriori (MAP) denoiser, motivating its replacement with a state-of-the-art image denoiser. While the interpretation as a minimization algorithm is no longer tenable under this substitution, it can be viewed as a fixed point algorithm that has been shown to converge under a variety of conditions [80, 89, 90].
PPP have been effectively adapted to solve CWI problems by integrating useful prior knowledge either through model-based methods or learning-based approaches. Aghamiry et al. [91] utilize PPP with the BM3D regularization algorithm and a proximal Newton method to address CWI ill-posedness challenges. However, their method struggles with the complexity of realistic noise. To address this limitation, learning-based methods have been developed that employ neural networks as denoising operators. For example, Fu et al. [92] show that deep CNNs can offer more effective regularization compared to traditional methods like total variation regularization. Similarly, Izzatullah et al. [93] design a CNN-based denoiser to sample from a regularized target posterior distribution, an approach refined to enhance uncertainty quantification in FWI.
@@ IV-C1 Uncertainty Quantification (UQ)
ML offers new opportunities to enable UQ for CWI by use of novel network architectures and approaches, such as normalizing flows[106, 107], deep probabilistic imaging [108], and generative AI (see next section), that can help to break the course of dimensionality and enable high-dimensional UQ. Here, we explore UQ estimation using neural networks, specifically through a method known as InvNet_UQ [109]. This approach focuses on aleatoric uncertainty and incorporates Simultaneous Quantile Regression (SQR) within a CNN framework [94]. InvNet_UQ employs a novel joint quantile regression loss function to enhance CWI by improving model reliability and quantifying prediction confidence. The network parameters of InvNet_UQ are obtained through the following optimization
@@
Liu et al. shows that data-driven models with UQ excel in handling contaminated data, including noisy or erroneous measurements. Their approach not only boosts confidence in inversion tasks but also refines training sets, enhancing model performance. Unlike Bayesian neural networks, InvNet_UQ does not necessarily require the image prior to be a Gaussian distribution, potentially providing more accurate uncertainty bounds compared to other methods. Beyond the use of quantile regression for UQ, Monte Carlo dropout—a technique favored for its simplicity—is also widely used in CWI. Zhu et al. [56] implement it by adding dropout layers to their network. Similarly, Schankee et al. [110] use a U-net architecture, employing both Monte Carlo dropout and bootstrap aggregating for uncertainty estimation. Yablokov et al. [111] use a multi-layer fully connected network and Monte Carlo simulations to estimate uncertainties.
@@ IV-C2 Generative AI (GenAI)
GenAI approaches can be broken down into two broad categories: sample-based approaches and score-based approaches. Sample-based approaches, such as Variational Auto-Encoders  [118] (VAEs) and Generative Adversarial Models [119] (GANs) learn the push forward from an easy-to-sample latent distribution (e.g., a multivariate i.i.d. Gaussian) to the data distribution. Once trained, they allow for direct (and computationally efficient) sampling. GANs, for example, were successfully employed to reduce the ill-posedness of CWI in the work of Zhang and Lin [71] and Mosser et al. [116] by leveraging the GAN latent space as a nonlinear dimension reduction technique, while VAEs were successfully employed in Yang et al. [64] for data augmentation. However, it is worth noticing that GAN and VAE approaches are generally limited to small/medium scale problems and may suffer from the so-called mode collapse phenomenon [120, 121, 122], which prevents them from fully capturing the distribution of the data. Score-based approaches overcome this by directly learning the Stein score of the data distribution (i.e. the gradient of the logarithm of the probability distribution function), which can be used in iterative sampling algorithms such as MCMC [123]. In particular, diffusion models [124, 125, 126] are now recognized as state-of-the-art in GenAI because of their high-quality sample generation, scalability to high-dimensional distributions, stability, and simplicity.
Wang et al. [114] use a diffusion model for solving CWI problems with quantified uncertainties. This method extends Song et al. ’s [126] approach for solving linear inverse problems (like CT and MRI image reconstruction) using diffusion models to the more challenging CWI problem, which is governed by a highly non-linear forward model. In this approach, a diffusion model is first constructed from velocity model data. Diffusion models consist of two components: a forward diffusion process that uses a stochastic ordinary differential equation (SDE) to map the data distribution to a fixed noise distribution, and a reverse diffusion process that maps the noise distribution to the data distribution. In particular, Wang et al. [114] adopt the variance preserving (VP) SDE [127] for the forward diffusion process, which can be written as
@@
[…]
@@
using white noise at the initial condition. To condition the sample generation to the measurement data, Wang et al. [114] include an FWI update after each step of discretized SDE solver to enhance data consistency.

### v2

@@ IV-B2 Physics-Aware Methods
Figure 8: Schematic illustration of a physics-informed neural network (PINN) [84], a physics-aware method for CWI that jointly learns the wavefield and velocity model by embedding the acoustic wave equation into the network’s loss function. Image courtesy of the authors of [84].
This category of methods integrates physical principles—most notably, the wave equation—directly into the learning framework through explicit constraints or regularization. A growing body of work has shown that enforcing the wave equation can significantly enhance inversion quality and physical interpretability in CWI [85, 11, 84, 86, 87, 88, 89].
A representative example [84] is the physics-informed neural network (PINN) approach [90], shown in Fig. 8. This model jointly performs wave propagation and inversion by embedding the governing wave equation into the loss function of a neural network. The architecture includes two neural networks: the first learns the wavefield $\phi(x,z,t)$ as a function of space and time, and the second learns the velocity model $\alpha(x,z)$, which serves as a key physical parameter in the wave equation.
The Physics-Informed Part imposes the 2D acoustic wave equation $\alpha^{2}\nabla^{2}\phi-\partial^{2}\phi/\partial t^{2}=0$, along with free-surface boundary conditions and snapshot constraints that align predicted wavefields with observed data at specific times. All differential operators are computed via automatic differentiation, allowing the model to enforce physical constraints without requiring meshing or numerical solvers. The total loss function combines the wave equation residual, boundary losses, and waveform mismatch, and is minimized via backpropagation until convergence. This mesh-free framework provides a flexible and differentiable solution for solving inverse problems, particularly in settings with limited prior information or complex subsurface structures.
Beyond PINNs, several physics-aware models adopt alternative network structures and optimization schemes. For instance, [85] introduces SWINet, a waveform inversion model that integrates wave equation solvers as differentiable layers and treats the velocity map as a trainable parameter. Similarly, [88] proposes FWIGAN, which fuses the wave equation with a Wasserstein GAN formulation to recover subsurface models in a generative framework. While these methods effectively embed physical constraints, they often rely on good initializations and face scalability challenges due to the high computational cost of solving wave equations, especially in 3D settings.
Physics-based regularization has also been incorporated into end-to-end learning pipelines. [11, 91, 86, 89] propose encoder–decoder networks that embed physical constraints to guide learning. [86], in particular, develops a framework that jointly performs forward modeling and inversion using an invertible neural network. These approaches strike a balance between physical interpretability and computational tractability, and are promising for tackling multiparameter and elastic wave imaging tasks.
@@ IV-B3 Parameterization Solution Methods
Figure 9: Schematic illustration of parameterization-based method [92]. Image courtesy of the authors of [92].
@@
Physics knowledge can be represented by the parameter space of the neural networks. [93] develops the deep image prior, a type of convolutional neural network, to learn useful prior knowledge via re-parameterization. A similar idea has been developed to solve CWI problems as shown in Fig. 9. In particular, velocity maps are parameterized while the forward model is used to calculate the associated waveform and obtain the data misfit. Mathematically, parameterization can be posed as
$\widehat{\bm{m}}$ $=g_{\bm{\theta}^{*}(\bm{d})}(\bm{\beta}),\;\mathrm{where}\;\bm{\theta}^{*}(\bm{d})=\operatorname*{arg\,min}_{\bm{\theta}}\mathcal{L}(f(g_{\bm{\theta}}(\bm{\beta})),\bm{d})\,$, (7)
where $\bm{\beta}$ is a random tensor drawn from a normal distribution; i.e., $\bm{\beta}\sim\mathcal{N}(0,\sigma^{2})$ and $g_{\bm{\theta}^{*}}(\cdot)$ is a network model used to parameterize the velocity maps. Unlike conventional physics-based methods that optimize velocity models directly in a high-dimensional voxel space, parameterization solution methods use a neural network to implicitly represent the velocity field. This has two key advantages. First, the network architecture acts as a strong implicit prior, promoting spatial smoothness, continuity, and structural plausibility—thereby regularizing the ill-posed inverse problem. Second, although the total number of network parameters can be large, the optimization is carried out over a lower-dimensional latent space (e.g., via $\bm{\beta}$ or a compact representation), and the network maps this to a structured velocity model. This effectively restricts the solution space to a subset of physically plausible models, improving convergence stability and robustness to noise.
@@
Parameterization solution methods, building on conventional physics-based approaches, inherit similar challenges. Typically, these methods require a well-defined initial velocity map, and factors like regularization and optimization techniques (e.g., gradients and step sizes) significantly influence their convergence. Specifically, crucial to Eq. (7) are the selection of $\bm{\beta}$ and the network structure as in $g_{\bm{\theta}^{*}}(\cdot)$. Early work by [94] samples $\bm{\beta}$ from noise and used CNNs. [67] adopts a similar approach but with a deeper theoretical focus. Both methods involve a two-step process: network pre-training with a specific initial guess, followed by fine-tuning. The discretization assumption limits the output to fixed sizes, restricting transferability. These limitations have motivated recent developments to improve flexibility and physical fidelity. Overcoming these drawbacks, [92] develops a new strategy that integrates the wave equation for regularization, eliminating the pre-training step and addressing the ill-posed nature of CWI problems. More recently, [66] introduces an approach that employs a continuous implicit function via neural networks to represent velocity maps, accommodating arbitrary sizes.
Figure 10: Schematic illustration of the network architecture of PPP-CWI [95], a Plug-and-Play Priors framework for CWI that alternates between physics-consistent updates and learned prior denoising. Image adapted from [95].
IV-B4 Plug and Play Priors Methods
The Plug-and-Play Priors (PPP) framework, initially developed in the computational imaging community [96, 97, 98], provides a flexible and powerful approach for solving inverse problems by integrating advanced image denoisers into iterative optimization algorithms. Originally designed to incorporate classical denoisers such as BM3D [99], PPP has since evolved to support modern learning-based models like DnCNN [100], enabling the use of expressive neural network priors without requiring problem-specific retraining.
PPP is rooted in proximal optimization techniques [101], particularly the Alternating Direction Method of Multipliers (ADMM) [102, 103, 104], which solves inverse problems of the form:
$\min_{\mathbf{m}}\;h(\mathbf{m})+r(\mathbf{m}),$
where $h(\cdot)$ typically represents a data fidelity term and $r(\cdot)$ serves as a regularization prior. The proximal operator, defined in [105, Ch. 6], is interpreted as a maximum a posteriori (MAP) estimator. This interpretation allows replacing it with advanced denoisers, including learned networks, while preserving algorithmic structure. Although this substitution breaks the strict optimization interpretation, it has been shown to converge under certain conditions [97, 106, 107].
In CWI, PPP has been adapted to handle the challenges of ill-posedness and data inconsistency. A representative framework, PPP-CWI [95], proceeds in two stages (shown in Fig. 10): an initial estimate is obtained from low-frequency data, followed by iterative refinement via a PPP-based scheme. The objective is:
$\min_{\bm{m}}\;g(\bm{m})+r(\bm{m}),\text{with }g(\mathbf{m})=\frac{1}{2}\|f(\bm{m})-\bm{d}\|_{2}^{2},$
where $f(\cdot)$ is the forward modeling operator and $\bm{d}$ denotes observed data.
The PPP-CWI algorithm iterates over three steps:
  1. 1.
    Model update (physics-consistent step):
    $\mathbf{m}^{(k+1)}=\arg\min_{\mathbf{m}}\;g(\mathbf{m})+\frac{\rho}{2}\left\|\mathbf{m}-\mathbf{y}^{(k)}+\mathbf{u}^{(k)}\right\|_{2}^{2},$ (8)
    typically solved using a first-order update:
    $\mathbf{m}^{(k+1)}=\mathbf{m}^{(k)}-\eta\left(\nabla g(\mathbf{m}^{(k)})+\rho(\mathbf{m}^{(k)}-\mathbf{y}^{(k)}+\mathbf{u}^{(k)})\right),$ (9)
    where $\eta$ is the learning rate and $\nabla g(\mathbf{m})$ is computed using the adjoint-state method.
  2. 2.
    Denoising update (prior-consistent step):
    $\mathbf{y}^{(k+1)}=\mathcal{D}_{\theta}\left(\mathbf{m}^{(k+1)}+\mathbf{u}^{(k)}\right),$ (10)
    where $\mathcal{D}_{\theta}$ is a pretrained denoising network (e.g., DRUNet [108] or DnCNN [100]).
  3. 3.
    Dual update:
    $\mathbf{u}^{(k+1)}=\mathbf{u}^{(k)}+\mathbf{m}^{(k+1)}-\mathbf{y}^{(k+1)}.$ (11)
This iterative scheme decouples the enforcement of physical consistency from the enforcement of prior knowledge, allowing both to be handled in a modular and flexible manner. Such decoupling is particularly valuable in CWI applications, where the forward model is highly complex, and the inversion problem is severely ill-posed. Traditional physics-based CWI methods often suffer from poor convergence due to non-convexity, high sensitivity to the initial model, and instability in the presence of noise or incomplete data. PPP mitigates these issues by incorporating powerful, data-driven priors, which is often learned from large datasets, directly into the inversion loop. These priors guide the solution toward physically plausible reconstructions, reduce reliance on handcrafted regularization terms, and improve robustness to data imperfections. Furthermore, by leveraging pretrained denoisers, PPP methods enhance generalization across different domains while reducing the need for extensive manual tuning. This makes them a promising and scalable alternative to conventional CWI, especially in large-scale or real-time imaging scenarios.
Several recent works have successfully applied PPP in CWI. [109] combines BM3D regularization with a proximal Newton solver to improve robustness in FWI, although performance degrades under complex noise. Learning-based alternatives have shown promise: [110] demonstrates that deep CNN priors outperform handcrafted regularization such as total variation, and [111] introduces a CNN-driven posterior sampling scheme to enhance uncertainty quantification in CWI.
IV-B5 Comparative Overview of Learning Strategies for CWI
This subsection provides a concise synthesis of the four learning strategies discussed in this section, namely, Simulations/Measurements-Driven, Physics-Aware, Parameterized Solution, and Plug-and-Play Priors. Each of these approaches brings distinct advantages and challenges. Summarizing their characteristics helps guide practitioners in selecting or combining appropriate strategies for specific CWI tasks.
Simulations/Measurements-Driven. These methods rely on supervised learning with paired data obtained from simulations or experimental measurements. Their primary strength lies in their ability to directly learn complex mappings from waveforms to physical properties when a large, diverse training dataset is available. However, their performance can degrade significantly under distribution shifts, particularly when real-world data differ from synthetic training conditions. These methods are most effective in well-controlled environments where high-quality paired data can be curated.
Physics-Aware. Physics-aware approaches incorporate physical knowledge, such as wave equations, either as soft constraints (e.g., through loss functions) or embedded in the architecture itself. By doing so, they reduce reliance on labeled data and improve generalization to unseen scenarios. However, these methods often involve more complex optimization procedures and may incur higher computational costs. They are particularly advantageous when physical models are well established, but labeled data are limited or expensive to obtain.
Parameterized Solution. In this strategy, a neural network is used to parameterize the velocity model, which is then optimized through a physics-based loss involving the forward model. The network acts as an implicit prior, constraining the inversion process to a lower-dimensional, structured space. This can improve convergence and regularization. Nonetheless, the approach is sensitive to the choice of network architecture, initialization, and training dynamics. It is well-suited for applications where capturing global structure or enforcing smoothness is critical.
Plug-and-Play Priors. Plug-and-play methods decouple the physics and data-driven components, allowing for the use of pretrained denoisers or generative models as implicit priors within iterative solvers. This modularity enables flexible integration and leverages powerful priors without altering the physics simulator. The trade-off, however, lies in computational overhead and the need to carefully balance the influence of priors and data consistency. These methods are promising in scenarios where data priors are rich, but direct supervision is limited.
Toward Strategy Selection and Integration. Each learning strategy offers a distinct balance among generalization ability, interpretability, data efficiency, and computational cost. Selecting an appropriate strategy depends not only on theoretical considerations but also on the practical constraints of the specific CWI application. For instance, in data-rich but poorly modeled environments, simulation-driven methods may suffice. Conversely, in data-sparse domains with reliable physics, physics-aware or parameterized methods are often more effective.
Recent trends suggest that hybrid approaches can leverage the strengths of multiple strategies. For example, plug-and-play frameworks can incorporate physics-aware denoisers or generative priors, such as score-based diffusion models, to improve inversion robustness without retraining the entire pipeline. Similarly, parameterized solvers can benefit from pretraining on synthetic data followed by fine-tuning with physics-constrained optimization, improving generalization while maintaining physical consistency.
Ultimately, strategy selection should consider factors such as the availability and quality of labeled data, the fidelity and computational cost of the forward model, the required level of interpretability, and tolerance to distribution shift. Designing adaptive systems that can switch between or combine strategies during training or inference remains an open but promising area for future research in CWI.
@@ IV-C1 Uncertainty Quantification (UQ)
ML offers new opportunities to enable UQ for CWI by use of novel network architectures and approaches, such as normalizing flows[124, 125], deep probabilistic imaging [126], and generative AI (see next section), that can help to break the curse of dimensionality and enable high-dimensional UQ. Here, we explore UQ estimation using neural networks, specifically through a method known as InvNet_UQ [127]. This approach focuses on aleatoric uncertainty and incorporates Simultaneous Quantile Regression (SQR) within a CNN framework [112]. InvNet_UQ employs a novel joint quantile regression loss function to enhance CWI by improving model reliability and quantifying prediction confidence. The network parameters of InvNet_UQ are obtained through the following optimization
@@
[127] shows that data-driven models with UQ excel in handling contaminated data, including noisy or erroneous measurements. Their approach not only boosts confidence in inversion tasks but also refines training sets, enhancing model performance. Unlike Bayesian neural networks, InvNet_UQ does not necessarily require the image prior to be a Gaussian distribution, potentially providing more accurate uncertainty bounds compared to other methods. Beyond the use of quantile regression for UQ, Monte Carlo dropout—a technique favored for its simplicity—is also widely used in CWI. [92] implements it by adding dropout layers to their network. Similarly, [128] uses a U-net architecture, employing both Monte Carlo dropout and bootstrap aggregating for uncertainty estimation. [129] uses a multi-layer fully connected network and Monte Carlo simulations to estimate uncertainties.
@@ IV-C2 Generative AI (GenAI)
GenAI approaches can be broken down into two broad categories: sample-based approaches and score-based approaches. Sample-based approaches, such as Variational Auto-Encoders  [136] (VAEs) and Generative Adversarial Models [137] (GANs) learn the push forward from an easy-to-sample latent distribution (e.g., a multivariate i.i.d. Gaussian) to the data distribution. Once trained, they allow for direct (and computationally efficient) sampling. GANs, for example, were successfully employed to reduce the ill-posedness of CWI in the work of [81] and [134] by leveraging the GAN latent space as a nonlinear dimension reduction technique, while VAEs were successfully employed in [74] for data augmentation. However, it is worth noticing that GAN and VAE approaches are generally limited to small/medium scale problems and may suffer from the so-called mode collapse phenomenon [138, 139, 140], which prevents them from fully capturing the distribution of the data. Score-based approaches overcome this by directly learning the Stein score of the data distribution (i.e., the gradient of the logarithm of the probability distribution function), which can be used in iterative sampling algorithms such as MCMC [141]. In particular, diffusion models [142, 143, 144] are now recognized as state-of-the-art in GenAI because of their high-quality sample generation, scalability to high-dimensional distributions, stability, and simplicity.
[132] uses a diffusion model for solving CWI problems with quantified uncertainties. This method extends [144] for solving linear inverse problems (like CT and MRI image reconstruction) using diffusion models to the more challenging CWI problem, which is governed by a highly non-linear forward model. In this approach, a diffusion model is first constructed from velocity model data. Diffusion models consist of two components: a forward diffusion process that uses a stochastic ordinary differential equation (SDE) to map the data distribution to a fixed noise distribution, and a reverse diffusion process that maps the noise distribution to the data distribution. In particular, [132] adopts the variance preserving (VP) SDE [145] for the forward diffusion process, which can be written as
@@
$\frac{1}{T}\int_{0}^{T}\lambda(t)\mathbb{E}_{\mathbf{m}_{t}|\mathbf{m}_{0}}\mathbb{E}_{\mathbf{m}_{0}}\left[\left\|s_{\theta}(\mathbf{m}_{t},t)+\frac{\mathbf{m}_{t}-e^{-\frac{1}{2}\int_{0}^{t}\beta(s)ds}\mathbf{m}_{0}}{1-e^{-\int_{0}^{t}\beta(s)ds}}\right\|_{2}^{2}\right]\,dt,$ (16)
@@
using white noise at the initial condition. To condition the sample generation to the measurement data, [132] include an FWI update after each step of discretized SDE solver to enhance data consistency.